OpenStudy (anonymous):

What is the solution to the equation? t(9 – 1) = 5 • 9 – 5 • 1

OpenStudy (anonymous):

t=5

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

here's the solution.. t(9-1)=5*9-5*1 9t-t=45-5 8t=40 t=5